Jumbo Juices preferred stock pays a constant dividend equal to $4.75 per share. The firms marginal tax rate is 40 percent.
Answer #1

Dividend = $ 4.75 per share

Marginal tax rate = 40%

Floatation cost = 5%

A. Firm issues 10,000 shares at $ 50 per share

Floatation cost = 5% of $ 50 = $ 2.5 per preferred stock

Net amount = 10,000 × (50 - 2.5)

= $ 475,000

B. Cost of preference share

\large r_{ps} = \frac{D_{ps}}{P(1-F)}

\large r_{ps} = \frac{4.75}{50(1-0.05)}

4.75 Tps 47.5

10% ps

Cost of preference share = 10%

Note in case of preference share there is no tax exemption on the dividend paid like debentures.

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me please help. Thank you.